OpenStudy (anonymous):

help please!!!!!:( Solve the equation for x. Round your answer to two decimal places. Do not include "x =" in your answer. e^(0.4x) = 0.4

OpenStudy (anonymous):

ln(0.4)/0.4

OpenStudy (anonymous):

is that the answer??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

im sorry i dont know how to do tht :/

OpenStudy (anonymous):

calculate ln value

OpenStudy (anonymous):

wat you mean nsutton?

OpenStudy (anonymous):

for approximate value, use calculator to approximate ln(0.4)

OpenStudy (anonymous):

so is -0.9163?

OpenStudy (anonymous):

-2.2907 after dividing by 0.4 is the final answer.

OpenStudy (anonymous):

so is -2.2907 divided by 0.4?

OpenStudy (anonymous):

Replace 0.4 with\[\frac{4}{10} \]and take the Log of each side. Then\[\frac{2 x}{5}=\text{Log}[2]-\text{Log}[5]\]

OpenStudy (anonymous):

so wats the answer?????:(

OpenStudy (anonymous):

\[\frac{5}{2} (\text{Log}[2]-\text{Log}[5])\text{ = }-2.2907268296853873 \]

OpenStudy (anonymous):

Sorry, missed the request for two decimal places. -2.29